# Dynamic LLM-Agent Network: An LLM-agent Collaboration Framework with Agent Team Optimization

### v1

@@ 5 Conclusion and Future Work
This work introduces a framework named Dynamic LLM-Agent Network (DyLAN) for LLM-agent collaboration on complicated tasks. DyLAN enables agents to interact for multiple rounds in a dynamic architecture with inference-time agent selection and an early-stopping mechanism to improve performance and efficiency. We further design an automatic agent team optimization algorithm based on an unsupervised metric termed Agent Importance Score, enabling the selection of best agents based on the contribution each agent makes. Overall, DyLAN reveals significant improvement on diverse tasks with relatively less computational cost compared to baselines.
In this work, the agents are driven by GPT-3.5-turbo and GPT-4, we will explore the effectiveness of DyLAN on agents built on open-source foundation models in future. And it is also worth to extend DyLAN to more complicate scenarios such as software development, virtual room chat, video games, and so on (Hong et al., 2023; Nascimento et al., 2023; Zhou et al., 2023; Zhu et al., 2023; Chan et al., 2023).
@@ Acknowledgments
Appendix A Broader Impact
LLM-agent systems are widely used in practical applications. DyLAN could also effortlessly cover practical software development, virtual room chat, video games, and so on (Hong et al., 2023; Nascimento et al., 2023; Zhou et al., 2023; Zhu et al., 2023; Chan et al., 2023). In these open-world environments, agents may operate as planners, actors, etc. DyLAN only requires people to give rough instructions on the constitute of agents and could automatically optimize a better team of agents to construct an efficient multi-agent system. These systems could benefit from DyLAN to reduce human labor on designing agents and have a better performance on their target tasks.
Also, the overall architecture of DyLAN (Figure 1) reflects the optimal collaboration organization of human online workers (Lykourentzou et al., 2022), and reveals significant performance in agent collaborations. Therefore, simulating human collaboration by LLM-agent collaborations under DyLAN might also be possible. Optimizing human collaboration by searching and simulating LLM agents will hopefully be more convenient and effective.
Appendix B Limitation
In experiments, we view code generation tasks as representative of open-ended generation tasks and adopt BLEU to decide whether two answers are consistent in Byzantine Consensus. In fact, the performance could be further leveraged by task-specific methods like CodeBLEU (Ren et al., 2020) or CodeT (Chen et al., 2023a).
For practical usage, the agent-evaluation metrics could cooperate with human annotation to give a more precise evaluation result, mainly when facing data scarcity problems. We apply agent team optimization on DyLAN simply on top of the inference-time agent selection. It still remains to be seen how to cooperate offline and online optimization methods in a finer granularity to further improve performance and efficiency in LLM-agent collaboration systems.
Appendix C Implementation Details
C.1 Detailed Experiment Settings
Data: Query $q$, DyLAN $S=({\mathbb{A}},\mathbb{E})$, task-specific answer extraction method $\mathrm{ans}$
Result: Output $o$
/* $\mathbb{E}=\{(a_{t,i},a_{t+1,j})\}_{t=1}^{T-1},a_{t,i},a_{t+1,j}\in{\mathbb{A}}$/
@@ C.1 Detailed Experiment Settings
       if $t=L$ then
             /* inference-time agent selection */
             ${\mathbb{R}}_{top}\leftarrow\mathrm{top-}m(\{r_{t-1,j}|a_{t-1,j}\in{\mathbb{A}}\})$;
             $\mathbb{E}\leftarrow\mathbb{E}\backslash\{(a_{t^{\prime},j},*),(*,a_{t^{\prime},j})|r_{t^{\prime},j}\in{\mathbb{R}}_{top},t^{\prime}\geq t-1\}$;
             $r_{t,j}\leftarrow r_{t-1,j},\forall(a_{t-1,j},a_{t,j})\in\mathbb{E}$;
@@
             $r_{t,i}\leftarrow a_{t,i}(p_{i},q,\{r_{t-1,j}|(a_{t-1,j},a_{t,i})\in\mathbb{E}\}),\forall i,\exists k,(a_{t,i},a_{t+1,k})\in\mathbb{E}$ ;
@@
      if Byzantine Consensus is reached then
             break;
@@
/* extract final answer */
$o\leftarrow\mathrm{ans}(\{r_{t,i}|a_{t,i}\in{\mathbb{A}}\},q)$;
Algorithm 1 The Inference Process $f_{\mathrm{Infer}}$ of DyLAN on an arbitrary query
Data: Output $o$, DyLAN $S=({\mathbb{A}},\mathbb{E})$
Result: Agent Importance Score of agents ${\bm{I}}$
$\mathrm{flag}\leftarrow\mathrm{False}$;
@@
                   $\mathrm{flag}\leftarrow\mathrm{True}$;
                   /* Initialzation */
                   distribute scores for $I_{t,i}$;
@@
                   ${\mathbb{R}}_{t-1}\leftarrow\{r_{t-1,j}|(a_{t-1,j},a_{t,i})\in\mathbb{E}\}$;
                   $[w_{t-1,1,i},...,w_{t-1,m,i}]\leftarrow a_{t,i}^{-1}(p_{i},q,{\mathbb{R}}_{t-1})$;
                   ${\bm{I}}_{t-1,j}\leftarrow{\bm{I}}_{t-1,j}+{\bm{I}}_{t,i}w_{t-1,j,i},a_{t-1,j}\in\{a_{t-1,j}|(a_{t-1,j},a_{t,i})\in\mathbb{E}\}$;
@@
Algorithm 2 The Calculation Process $f_{\mathrm{Imp}}$ of Agent Importance Score within DyLAN
@@
In all experiments, we use gpt-35-tubo-0301 for every LLM agent if not specified. In Table 4, “(GPT-4)” denotes gpt-4-0613 and “(Codex)” denotes code-davinci-002 from OpenAI (Chen et al., 2021; OpenAI, 2023). To avoid the context length issue in prior work (Du et al., 2023; Liu et al., 2023a), we set memory space for agents in DyLAN to $1$ only to keep the freshest responses of predecessors. We set max tokens to 2048 to avoid exceeding the maximum context length. We use a listwise ranker in the inference-time agent selection of DyLAN because of the effectiveness and efficiency, compared to ELo rating (Herbrich et al., 2006) or Sliding Window (Qin et al., 2023) we have tested in Appendix D.3. We use the same ranker to implement LLM-Blender (Jiang et al., 2023) in experiments. We set $m=k=2$ in the inference-time agent selection and agent team optimization. To avoid positional bias, we shuffle the responses from agents at $(L-1)$-th time step before performing inference-time agent selection. The detailed algorithm is in Algorithm 1. To implement the early-stopping mechanism, we need to determine whether the answers from the nodes in the same layer of DyLAN are consistent. For classification problems, the answers are consistent if identical, and for open-ended generation, the consistency is determined by a threshold of BLEU score.
@@
In general reasoning, we extract the answer from the response by matching the last “(X” or “(X)”, where “X” represents A, B, C or D. On average, DyLAN with the 7 agents consumes 8.30 API calls during the agent team optimization. Inference-time agent selection functions on the third round ($L=3$ in Algorithm 1). They could go through at maximum $T=4$ rounds of interaction. We also searched temperature in {0, 0.2, 0.8, 1.0} for the best configuration for each system. In arithmetic reasoning, we set temperature to 0 for the single execution and PHP, 0.2 for LLM Debate, LLM-Blender, and DyLAN with Complex CoT prompts, and 1.0 for DyLAN with simple CoT prompts in Table 2, since systems with the same prompts will give all the same responses if temperature is zero, causing degradation. Also, we did not optimize the team of agents for DyLAN on arithmetic reasoning tasks. We follow the answer extraction method from the origin paper (Hendrycks et al., 2021b). We construct DyLAN with 4 agents assigned no specific roles and let agents to interact for at maximum 4 rounds. We reported the numbers of API calls of running DyLAN on the optimized team of agents. In DyLAN, inference-time agent selection functions at the third time step ($L=3$). Furthermore, we use the exact match to determine the consistency of answers in the early-stopping mechanism and extract the final answer from nodes in the last layer on reasoning tasks.
@@
In the code generation task, we set temperature to 0 for the single execution, Reflexion, and DyLAN, and 0.8 for LLM Debate, LLM-Blender, CodeT, and DyLAN in Table 4. In DyLAN, we optimized four agents to write code and four agents to give feedback from the last 12 candidate roles in Appendix E. The optimization is also conducted on gpt-35-tubo-0301. The selected code writers are “Python Assistant”, “Algorithm Developer”, “Computer Scientist”, and “Programmer”; and the selected judges are “Syntax Checker”, “Unit Tester”, “Reflector”, and “Ranker”. “Syntax Checker” is pure external tools without LLMs and “Unit Tester” is equipped with a code interpreter. We reported the numbers of API calls of running DyLAN on the optimized team of agents. On average, during the agent team optimization, DyLAN with the 12 agents consumes 23.04 API calls. In DyLAN, solutions given by code writers are reviewed by judges in at maximum $T=6$ rounds. At time step $t=1,3,4,6$, code writers gives solutions and judges review it at $t=2,5$ ($L=4$). Specifically, early-stopping mechanism functions at the third layer and later ($t\geq 3$). We use BLEU score in the early-stopping mechanism. We calculate BLEU by sacreBLEU [^2] The signature of sacreBLEU is “nrefs:1|normal-|||case:mixed|normal-|||eff:no|normal-|||tok:13a|normal-|||smooth:exp|normal-|||version:2.3.1”. (Post, 2018). For answer extraction, we store all unit tests from the unit tester (if exists in the system) and randomly select the final output from the top 5 code completions from all nodes that pass most tests.
C.2 Calculation of Agent Importance Score
To implement the agent team optimization algorithm under DyLAN, only one sentence needs to be injected into the end of the prompt of each node in DyLAN: Along with the answer, give a score ranging from 1 to 5 to the solutions of other agents. Put all $\{num_{\mathrm{predecessor}}\}$ scores in the form like [[1, 5, 2, ...]], where $num_{\mathrm{predecessor}}$ denotes the number of predecessors of the node. The prompt functions as the $f_{t,i}^{(s)}$ in Section 3.1 and we extract $w_{t,j,i}$ from its response at the same time when we extract the message that passes between nodes. To avoid positional bias, we shuffle responses from agents at previous time step when rating.
In Algorithm 2, initial contributions are distributed on nodes at the last layer. For reasoning tasks, we uniformly distribute contributions to agents that give consistent answers in the last layer. On code generation tasks, we uniformly distribute contributions in the final round with no syntax error in their answers. During agent team optimization, we independently optimize the composition of agents for each domain. Thus, in experiments on general reasoning and code generation tasks, we sum up agent importance scores to optimize a team of agents for each subject on the MMLU dataset and for the HumanEval dataset, respectively.
Appendix D Additional Results
@@
D.1 Human Priors and Automatic Evaluation Results
Table 7: Subjects on which agents have the top-ranked Agent Importance Score in the experiment with DyLAN of 7 agents on MMLU dataset. Green annotation denotes the fields related to the role from the human perspective, which are annotated manually.

### v2

@@ 5 Conclusion and Future Work
This work introduces a framework named Dynamic LLM-Powered Agent Network (DyLAN) for collaboration of dynamic agent teams on complicated tasks. DyLAN functions in a two-stage paradigm, enabling agents to interact in a dynamic structure with agent team reformation. In “Team Optimization” stage, the agent selection algorithm based on an unsupervised metric termed Agent Importance Score, selects top contributory agents in a principled way for collaboration on “Task Solving”. Overall, DyLAN reveals improvement on diverse tasks with relatively less computational cost compared to baselines. In the future, we plan to explore the effectiveness of DyLAN built on open-source foundation models.
@@ Acknowledgments
Ethics Statement
LLM-powered agent systems are widely used in practical applications. DyLAN could also effortlessly cover practical software development, virtual room chat, video games, and so on (Hong et al., 2024; Nascimento et al., 2023; Zhou et al., 2023; Zhu et al., 2023; Chan et al., 2024). In these open-world environments, agents may operate as planners, actors, etc. DyLAN only requires people to give rough instructions on the constitute of agents and could automatically optimize a better team of agents to construct an efficient multi-agent system. These systems could benefit from DyLAN to reduce human labor on designing agents and have a better performance on their target tasks.
Also, the overall architecture of DyLAN (Figure 2) reflects the optimal collaboration organization of human online workers (Lykourentzou et al., 2022), and reveals significant performance in agent collaborations. Therefore, simulating human collaboration by LLM-powered agent collaborations under DyLAN might also be possible. Optimizing human collaboration by searching and simulating LLM agents will hopefully be more convenient and effective. We also acknowledge the potential risk in pretrained language models used in the paper, e.g., GPT-3.5 and GPT-4, which may cause improper responses. Furthermore, creating agents by hand and LLM generations might prompt LLM-powered agents to act or response misaligned with principles of society, which may possibly happen during agent collaborations. However, we think team optimization process could potentially alleviate this situation.
Appendix A Discussion & Limitation
In experiments, we view code generation tasks as representative of open-ended generation tasks and adopt BLEU to decide whether two answers are consistent in early stopping mechanism in Section 3.3.2. In fact, the performance could be further leveraged by task-specific methods like CodeBLEU (Ren et al., 2020) or CodeT (Chen et al., 2023a).
For practical usage, the agent-evaluation metrics could cooperate with human annotation to give a more precise evaluation result on individual contributions of agents, mainly when facing data scarcity problems. Furthermore, we simply incorporating agent selection on DyLAN with agent team reformation, as a primary step towards collaboration of dynamic agent teams. It still remains to be seen how to cooperate off-collaboration and in-collaboration optimization methods in a finer granularity to further improve performance and efficiency in LLM-powered agent collaboration systems.
Additionally, though agent selection could differentiate top contributory agents, in extreme cases where the majority of agents are designed to contradict the task requirement, low performance might be caused, e.g., agents are prompted or trained to generate codes may face difficulties in clinical question answering. To tackle the imbalance of high- and low-performing agents, replicating agents with high Agent Importance Score instead of including low-score agents could be a solution. Additionally, in extreme circumstances, we can automatically introduce agents from more capable LLMs with validation, in addition to agent selection.
Appendix B Implementation Details
B.1 Detailed Experiment Settings
Algorithm 1 The Inference Process of DyLAN on an Arbitrary Query
  Input: T-FFN $\mathcal{G}=(\mathcal{V}_{1},\cdots,\mathcal{V}_{T};E_{1},\cdots,E_{T-1,T})$, Query $q$
  Output: Final Answer $o$
  // $E=\{(v_{t,i},v_{t+1,j})\}_{t=1}^{T-1},v_{t,i},v_{t+1,j}\in\mathcal{V}$
  // $\phantom{E}=\bigcup_{t=1}^{T}\mathcal{V}_{t}$.
  // $m_{t,i}\in\mathcal{M}_{t}$ denotes the response from $v_{t,i}\in\mathcal{V}_{t}$.
@@ B.1 Detailed Experiment Settings
     if agent team reformation at time step $t$ then
        $\mathcal{M}_{top}\leftarrow\mathrm{top-}k(\{m_{t-1,j}|v_{t-1,j}\in\mathcal{V}_{t-1}\})$
        $E\leftarrow E\backslash\{(v_{t^{\prime},j},v_{t^{\prime\prime},j^{\prime}}),(v_{t^{\prime\prime},j^{\prime}},v_{t^{\prime},j})|$
        $\quad m_{t^{\prime},j}\in\mathcal{M}_{top},m_{t^{\prime},j^{\prime}}\notin\mathcal{M}_{top},t^{\prime},t^{\prime\prime}\geq t-1\}$
        $m_{t,j}\leftarrow m_{t-1,j},\forall(v_{t-1,j},v_{t,j})\in E$
@@
        $\forall i,\exists k,(v_{t,i},v_{t+1,k})\in E_{t,t+1}$,
        $m_{t,i}\leftarrow$
        $\quad f_{\mathrm{mp}}(\{m_{t-1,j}|(v_{t-1,j},v_{t,i})\in E_{t,t+1}\},v_{t,i})$
@@
     if early stopping then
        $T\leftarrow t$
        break
@@
  $o\leftarrow\mathrm{postProcess}(\mathrm{maxCount}\{\mathcal{M}_{T}\})$
Algorithm 2 The Team Optimization Process $f_{\mathrm{Optim}}$ of Agent Importance Score within DyLAN
  Input: Output $o$, T-FFN $\mathcal{G}=(\mathcal{V},E)$
  Output: Agent Importance Score of agents $\bm{I}$
  // $m_{t,i}\in\mathcal{M}_{t}$ denotes the response from $v_{t,i}\in\mathcal{V}_{t}$.
  $\mathrm{flag}\leftarrow\mathrm{False}$
@@
           $\mathrm{flag}\leftarrow\mathrm{True}$
           distribute scores for $I_{t,i}$
@@
           $\mathcal{M}_{t-1}\leftarrow$
           $\quad\{m_{t-1,j}|(v_{t-1,j},v_{t,i})\in E\}$
           $[w_{t-1,1,i},...,w_{t-1,m,i}]\leftarrow$
           $\quad f_{t,i}^{(s)}(p_{i},q,\mathcal{M}_{t-1})$
           $\bm{I}_{t-1,j}\leftarrow\bm{I}_{t-1,j}+\bm{I}_{t,i}w_{t-1,j,i},$
           $\quad\exists(v_{t-1,j},v_{t,i})\in E$
@@
In all experiments, we use gpt-35-tubo-0301 for every method if not specified. The version of GPT-4 is GPT-4-0613. In Table 2, “(Codex)” denotes code-davinci-002 from OpenAI (Chen et al., 2021; OpenAI, 2023). All experiments with non-zero temperature is repeated for three times and the median is reported. To avoid the context length issue in prior work (Du et al., 2023; Liu et al., 2024), we set memory space for agents in DyLAN to $1$ only to keep the freshest responses of predecessors. We set max tokens to 2048 for GR and AR tasks and 1024 for CG and DM tasks to avoid exceeding the maximum context length. The construction of candidates are demonstrated in Appendix D. We set $N=4$ in T-FFN after team optimization because the early-stopping mechanism requires at least four agents to tolerate one different response at a specific time step (Section 3.3.2); when reaching consensus over 2/3 of agents, it allows for $14-\left\lceil{\frac{2}{3}N}\right\rceil=$ 1 excetional response. We use a listwise ranker in the agent team reformation of DyLAN because of the effectiveness and efficiency, compared to ELo rating (Herbrich et al., 2006) or Sliding Window (Qin et al., 2023) we have tested in Section C.5. We use the same ranker to implement LLM-Blender (Jiang et al., 2023) in experiments. We set $k=2$ in the agent team reformation, because it’s the minimal number for collaborations and we empirically found it brings great trade-off between effectiveness and efficiency. To avoid positional bias, for each time step $t$, we shuffle the responses from agents at $t-1$ when passing messages towards agents at $t$. The detailed inference algorithm is in Algorithm 1. To implement the early-stopping mechanism, we need to determine whether the answers from the nodes in the same layer of DyLAN are consistent. For classification and decision-making problems, the answers are consistent if identical, and for open-ended generation, the consistency is determined by a threshold of BLEU score.
@@
In general reasoning, we extract the answer from the response by matching the last “(X” or “(X)”, where “X” represents A, B, C or D. On average, Agent team reformation functions on the third time step. They could go through at maximum $T=4$ rounds of interaction. We also searched temperature in {0, 0.2, 0.8, 1.0} for the best configuration for each system. In arithmetic reasoning, we set temperature to 0 for the single execution and PHP, 0.2 for LLM Debate, LLM-Blender, and DyLAN with Complex CoT prompts, and 1.0 for DyLAN with simple CoT prompts in Table 3, since systems with the same prompts will give all the same responses if temperature is zero, causing degradation. Prompting templates are replicated from their original studies, including normal CoT prompts (Wei et al., 2022) from the MATH dataset (Hendrycks et al., 2021b) and Complex CoT from PHP (Zheng et al., 2023). We follow the answer extraction method from the origin paper (Hendrycks et al., 2021b). We construct DyLAN with 4 agents assigned no specific roles and let agents to interact for at maximum $T=4$ rounds under T-FFN formulation. We reported the classification accuracy of each category averaged across subjects and the numbers of API calls of running DyLAN on the optimized team of agents.
@@
[…]
Experiments on Decision Making Tasks
In the decision-making task, we set temperature to 0 for DyLAN and all baselines in Table 2. For ReAct with self-consistency (denoted by ReAct-SC in the table) (Wang et al., 2023b), we sampled three times for each response. In DyLAN, we optimized four agents from 8 candidates which are depicted in Appendix D. All methods are also conducted on gpt-35-tubo-1106. We did not select LASER (Ma et al., 2023) as a baseline, because it requires GPT-4 for better performance and it extracts all valid actions in each page into function calls, instead of detected by agent itself, which we decide to be a different setting. We divide the pages of the WebShop environment into 3 parts: the initialization page for “searching” part, the item list page for “exploring” part, and the item details pages for “item” part. Thus, we managed to optimize teams for each part from agents in Appendix D: “Search Optimizer”, ’‘Budget Analyst”, “Instruction Analyst”, “Decision Reflector” for “searching” group, “Decision Maker”, “Budget Analyst”, “Product Explorer”, “InstructionAnalyst” for “exploring” group, and “Budget Analyst”, “Description Reader”, “Decision Maker”, “Result Estimater” for “item” group. Agents interact for at maximum $T=4$ steps for each action. We simply concatenate observations of previous actions on each decision. For answer post-processing, we skip invalid actions from the outputs of $V_{T}$.
B.2 Calculation of Agent Importance Score
To implement the agent selection algorithm under DyLAN, only one sentence needs to be injected into the end of the prompt of each node in T-FFN: “Along with the answer, give a score ranging from 1 to 5 to the solutions of other agents. Put all $\{num_{\mathrm{p}}\}$ scores in the form like [[1, 5, 2, ...]]”, where $num_{\mathrm{p}}$ denotes the number of predecessors of the node. The prompt functions as the $f_{t,i}^{(s)}$ in Section 3.4 and we extract $w_{t,i,j}$ from its response at the same time when we extract the message that passes between nodes. The scores are normalized so that their sum ($\sum_{i=1}^{N}w_{t,i,j}$) equals 1. To avoid positional bias, responses from agents at previous time step are shuffled when rating.
In Algorithm 2, initial contributions are distributed on nodes at the last layer. For reasoning and dicision-making tasks, we uniformly distribute contributions to agents that give consistent answers in the last layer. On code generation tasks, we uniformly distribute contributions in the final round with no syntax error in their answers.
Appendix C Additional Results
@@
Indicator Dataset GR CG
NA - 63.5 76.2
Random - 64.8 75.6
Human Prior - 66.7 78.0
Agent 1% 68.9 79.3
Imp. Score 10% 72.2 82.3
100% 73.6 82.9
Table 9: Experimental results of different indicators used in agent selection during team optimization in DyLAN on five subjects in the GR and CG tasks. “Dataset” denotes the proportion of dataset used in team optimization.
C.1 Data Efficiency of Team Optimization
We further demonstrate the data efficiency of agent selection by performing it based on different amounts of data. The experiments are conducted on five subjects in the GR task (the same as Table 8) and the CG task. We sample the subsets with the proportions of 1% and 10% of the original dataset. Agent Importance Score for agent selection is averaged on the subsets, and the selected team is tested on the whole dataset. We raise random selection and human prior selection as baselines. The latter is simulated by GPT-4 prompted by the task and agent descriptions (Appendix D). As shown in Table 9, by optimizing the team 10% of the original dataset, DyLAN has demonstrated similar performance compared to using the whole dataset, with only 0.2 loss on GR and 0.6 loss on CG. We can observe that even with only 1% of the original dataset, DyLAN could obtain a significant improvement of +3.7 over random selection on CG. From observation, agents augmented with tools are always selected during team optimization under different proportions of the dataset, indicating the effectiveness of Agent Importance Score as an indicator. Please refer to Section C.3 for a detailed analysis of the human priors.
C.2 Robustness of different foundation models in DyLAN
Method Pass@1 #API Calls
Single-Agent Methods
Single Execution 88.4 (+0.0) 1.00
LATS 94.4 (+6.0) $>$40.00
Reflexion 91.4 (+3.0) 7.32
Multi-Agent Methods
Meta-GPT 85.9 (-3.5) $>$30.00
AgentVerse 89.0 (+0.6) 27.00
DyLAN (Ours) 92.1 (+3.7) 15.94
Table 10: Experimental results on the CG task on GPT-4-0613. The number in parentheses indicates the difference relative to the single execution or direct execution. The bold font denotes the results of our method and the best results are underlined.
Besides using GPT-3.5 for DyLAN on CG tasks in Table 2, we also experiment with GPT-4 in Table 10. Due to budget limits, we directly reuse the performance reported in the paper of baselines, including LATS (Zhou et al., 2023), Reflexion (Shinn et al., 2023), Meta-GPT (Hong et al., 2024), and AgentVerse (Chen et al., 2024), and estimate the cost in terms of numbers of API calls. DyLAN is also constructed by agents which are optimized based on GPT-3.5, as demonstrated in Section B.1. We found that DyLAN consistently outperforms other multi-agent methods, indicating the effectiveness of dynamic agent team in T-FFN structure and the cross-model transferability of optimization results on agent teams. Although LATS outperforms DyLAN, it requires over 40 times GPT-4 calls per sample to conduct inference-time MCTS on GPT-4, which demonstrates poor efficiency.
C.3 Human Priors and Agent Importance Scores
We further investigated how these agents selected by our unsupervised metric Agent Importance Score differ from human priors (e.g., these predefined roles). To do so, we calculated agent importance scores for 7 agents on each subject of the MMLU dataset. As an example, we show the subjects where the agent of “Doctor” and “Programmer” has the highest agent importance score among all agents in Table 11 and Table 12.